# Hierarchical Indexing for Retrieval-Augmented Opinion Summarization

### v1

@@ 6.2 Automatic Evaluation
We evaluate summary generation using Space (Section 4.2), which includes multiple reference summaries created by human annotators for each entity. We also include AmaSum Bražinskas et al. (2021), to evaluate summary generation on a wide range of categories of Amazon products, with an average of 560 reviews per entity. The reference summaries were collected from professional review websites, and therefore are not necessarily grounded in the input reviews. We use the same splits, based on four product categories, as released by Hosking et al. (2023b). Further dataset statistics are given in LABEL:appx:datasets.
@@
We select a random review from the inputs as a lower bound. We include an extractive oracle as an upper bound, by selecting the input sentence with highest ROUGE-2 similarity to each reference sentence. [^4] When multiple references are available, we select one at random. For a k-means baseline, we run $k$-means on MiniLM sentence embeddings Reimers and Gurevych (2019), then extract the nearest sentence to the cluster centroids. We set $k=8$ to match the average sentence length of the reference summaries. Lexrank Erkan and Radev (2004) is an unsupervised extractive method using graph-based centrality scoring of sentences. QT Angelidis et al. (2021) uses vector quantization to map sentences to a discrete encoding space, then generates extractive summaries by selecting representative sentences from clusters. SemAE Basu Roy Chowdhury et al. (2022) is an extractive method that extends QT, relaxing the discretization and encoding sentences as mixtures of learned embeddings. CopyCat Bražinskas et al. (2020) is an abstractive approach that models sentences as observations of latent variables representing entity opinions, trained in a ‘leave one out’ manner. BiMeanVAE and COOP Iso et al. (2021) are abstractive methods that encode full reviews as continuous latent vectors using an autoencoder, and take the average (BiMeanVAE) or an optimised combination (COOP) of review encodings. We compare to a recent open weight instruction-tuned LLM, specifically Mistral Instruct v0.2 7B Jiang et al. (2023). To evaluate the effect of model size, we compare to Llama 2 Chat 7B and 13B Touvron et al. (2023), which are equivalent models apart from parameter count. The LLMs are prompted zero-shot using the prompt in Appendix C, and sampled with temperature $0.7$. We report the mean and standard deviation scores based on 3 samples.
Most of the abstractive methods are not scalable and have upper limits on the number of input reviews. CopyCat and the LLMs have a maximum input sequence length, while COOP exhaustively searches over combinations of input reviews. We use 8 randomly selected reviews as input to CopyCat, COOP, Llama 2, and Mistral 7B.
@@
The results in Table 2 show that HIRO achieves the highest SAP scores across both datasets, indicating that it generates summaries with the best balance between absolute prevalence and specificity. While CopyCat and COOP achieve the highest prevalence scores on Space and AmaSum respectively, they also display some of the highest genericness; qualitatively, the outputs are very similar to each other, and include no specific details about each entity. We give example outputs in Tables LABEL:tab:output and LABEL:tab:output2, and an example of selected subpaths, sentence clusters and corresponding HIRO outputs in LABEL:tab:evidence_example.
Increasing model size does not help
The SAP scores decrease for the larger Llama 2 13B model compared to Llama 2 7B, indicating that increasing model size does not lead to improved performance. The short context window of current instruction-tuned LLMs limits the number of reviews they can take as input, preventing them from accurately reflecting the true distribution of opinions.
@@
Figure 4: Distribution of selected subpaths by depth. A significant fraction of the extracted clusters come from paths deeper than the top level.
@@
Figure 5: Results of our human evaluation. Crowdworkers were asked for pairwise preferences between generated summaries in terms of their Accuracy, Detail, and Coherence & Fluency, as well as an overall preference. HIRO generates more coherent and detailed summaries that better represent the opinions in the input reviews than comparison systems, and are preferred by human annotators.
HIRO is more attributable
To evaluate how faithful the generated summaries are to the retrieved sentence clusters or evidence sets, we use an NLI model to determine how many sentences in each cluster either entail or are entailed by the corresponding sentence in the output summary, and take the mean. Considering both forward and backward entailment in this manner accounts for the different levels of granularity between the inputs and summary Zhang et al. (2024); input reviews are likely to be more specific than summary sentences, but concise summary sentences are likely to contain multiple assertions, e.g. “The food was good and the rooms were clean”. HIROdoc does not align the evidence sets with each generated sentence, so we calculate the maximum support score over all sets for each summary sentence. Most abstractive systems are not attributable, and so we only compare with Hercules. Table 3 shows the percentage of summary sentences that have support from at least one sentence in the evidence (partial support) and from at least half the sentences in the evidence (majority support). HIROdoc generates summaries that have strong partial support from the associated evidence sets, with improved majority support compared to Hercules despite also being significantly more detailed.
@@ 6.3 Human Evaluation
We conduct a human evaluation to verify that HIRO generates summaries that are coherent and accurately reflect the opinions in the input reviews. We recruit crowdworkers through Prolific, and show them a set of 50 reviews, followed by two generated summaries. We solicit pairwise preferences along three dimensions, as well as an overall preference:
@@
The full instructions are reproduced in Appendix A. We gather annotations for 10 entities each from the Space and AmaSum test sets, with 3 annotations for each.
@@
The results in Figure 5 show that HIROdoc produces summaries that outperform comparison systems across all dimensions, producing summaries that coherently and accurately represent the opinions in the input reviews. Differences between HIROdoc and other systems are significant in all cases (using a one-way ANOVA with post-hoc Tukey HSD test, $p<0.05$), except for coherence versus Mistral 7B. Both Mistral 7B and HIROdoc outperform the reference summaries, supporting findings from prior work Bhaskar et al. (2023); Hosking et al. (2023a).
Figure 6: Results of a human evaluation comparing the three variants of HIRO: extractive, sentence-wise and document. Overall, annotators prefer the coherent of the document approach, but the sentence-wise variant generates more detailed summaries that are also more attributable. The preferred tradeoff between coherence and attribution will vary depending on the application.
@@ 7 Conclusion
We propose HIRO, a modular method for unsupervise opinion summarization that uses a hierarchical index over sentences to select clusters of prevalent opinions. Our approach leverages pretrained Large Language Models to generate coherent and fluent summaries that are attributable and accurately reflect the popularity of opinions in the input reviews. Extensive experiments show that, as well as generating higher quality summaries, HIRO learns a more semantically distributed representation than competitive baselines. While we limit our experiments to opinion summarization, we believe that HIRO could be usefully applied to a wide range of other retrieval-augmented generation tasks.
@@ Appendix B Hyperparameters
 Optimizer Adam Kingma and Ba (2015)
@@ Appendix C LLM prompts
Prompt 1: Baseline LLMs

### v2

@@ 6.2 Automatic Evaluation
We evaluate summary generation using Space (Section 4.2), which includes multiple reference summaries created by human annotators for each entity. We also include AmaSum (Bražinskas et al., 2021), to evaluate summary generation on a wide range of categories of Amazon products, with an average of 560 reviews per entity. The reference summaries were collected from professional review websites, and therefore are not necessarily grounded in the input reviews. We use the same splits, based on four product categories, as released by Hosking et al. (2023b). Further dataset statistics are given in LABEL:appx:datasets.
@@
We select a random review from the inputs as a lower bound. We include an extractive oracle as an upper bound, by selecting the input sentence with highest ROUGE-2 similarity to each reference sentence. [^4] When multiple references are available, we select one at random. For a k-means baseline, we run $k$-means on MiniLM sentence embeddings (Reimers and Gurevych, 2019), then extract the nearest sentence to the cluster centroids. We set $k=8$ to match the average sentence length of the reference summaries. Lexrank (Erkan and Radev, 2004) is an unsupervised extractive method using graph-based centrality scoring of sentences. QT (Angelidis et al., 2021) uses vector quantization to map sentences to a discrete encoding space, then generates extractive summaries by selecting representative sentences from clusters. SemAE (Basu Roy Chowdhury et al., 2022) is an extractive method that extends QT, relaxing the discretization and encoding sentences as mixtures of learned embeddings. CopyCat (Bražinskas et al., 2020) is an abstractive approach that models sentences as observations of latent variables representing entity opinions, trained in a ‘leave one out’ manner. BiMeanVAE and COOP (Iso et al., 2021) are abstractive methods that encode full reviews as continuous latent vectors using an autoencoder, and take the average (BiMeanVAE) or an optimised combination (COOP) of review encodings. We compare to a recent open weight instruction-tuned LLM, specifically Mistral Instruct v0.2 7B (Jiang et al., 2023). Since no training data is available, the LLM was prompted zero-shot as per Appendix C, and sampled with temperature $0.7$. We report the mean and standard deviation scores based on 3 samples.
Most of the abstractive methods are not scalable and have upper limits on the number of input reviews. CopyCat and Mistral 7B have a maximum input sequence length, while COOP exhaustively searches over combinations of input reviews. We use 8 randomly selected reviews as input to CopyCat, COOP, and Mistral 7B.
@@
The results in Table 2 show that HIRO achieves the highest SAP scores across both datasets, indicating that it generates summaries with the best balance between absolute prevalence and specificity. While CopyCat and COOP achieve the highest prevalence scores on Space and AmaSum respectively, they also display some of the highest genericness; qualitatively, the outputs are very similar to each other, with few specific details. We give example output in Tables LABEL:tab:output and LABEL:tab:output2.
@@
Figure 4: Distribution of selected subpaths by depth. A large fraction of the extracted clusters come from paths deeper than the top level.
Figure 5: Overall results of our human evaluation based on the test splits of Space and AmaSum. Crowdworkers were asked for pairwise preferences between generated summaries in terms of their Accuracy, Detail, and Coherence & Fluency, as well as an overall preference. HIRO generates more coherent and detailed summaries that better represent the opinions in the input reviews than comparison systems, and are preferred by human annotators.
HIRO is more faithful to selected evidence
To evaluate how faithful the generated summaries are to the retrieved sentence clusters or evidence sets, we use an NLI model to determine how many sentences in each cluster either entail or are entailed by the corresponding sentence in the output summary, and take the mean. Considering both forward and backward entailment in this manner accounts for the different levels of granularity between the inputs and summary (Zhang et al., 2024); input reviews are likely to be more specific than summary sentences, but concise summary sentences are likely to contain multiple assertions, e.g. “The food was good and the rooms were clean”. HIROdoc does not align the evidence sets with each generated sentence, so we calculate the maximum support score over all sets for each summary sentence. Most abstractive systems are not attributable, and so we only compare with Hercules. Table 3 shows the percentage of summary sentences that have support from at least one sentence in the evidence (partial support) and from at least half the sentences in the evidence (majority support). HIROdoc generates summaries that have strong partial support from the associated evidence sets, with improved majority support compared to Hercules despite also being significantly more detailed.
@@ 6.3 Human Evaluation
We conduct a human evaluation to verify that HIRO generates summaries that are coherent and accurately reflect the opinions in the input reviews. We recruit crowdworkers through Prolific, selected to be L1 English speakers from the US or UK with a minimum of 100 previously approved studies, compensated above the UK living wage at 12GBP/hr. Participants were allowed to rate at most 5 samples. We show annotators a set of 50 reviews (chosen based on pilot studies to balance annotator load with reliability), followed by two generated summaries. We solicit pairwise preferences (Louviere and Woodworth, 1990) along three dimensions, as well as an overall preference:
@@
The full instructions are reproduced in Appendix A. Ties (i.e., ‘no difference’) were allowed. We gather annotations for 10 entities each from the Space and AmaSum test sets, with 3 annotations for each pairwise combination of system outputs, leading to a total of 900 pairwise ratings. The study was approved by ethics committee, ref. #491139.
@@
The results in Figure 5 show that HIROdoc produces summaries that outperform comparison systems across all dimensions, producing summaries that coherently and accurately represent the opinions in the input reviews. Differences between HIROdoc and other systems are significant in all cases (using a one-way ANOVA with post-hoc Tukey HSD test, $p<0.05$), except for coherence versus Mistral 7B. Both Mistral 7B and HIROdoc outperform the reference summaries, supporting findings from prior work (Bhaskar et al., 2023; Hosking et al., 2023a).
Figure 6: Results of a human evaluation comparing the three variants of HIRO: extractive, sentence-wise and document. Overall, annotators prefer the coherence of the document approach, but the sentence-wise variant generates more detailed summaries that are also more attributable. The preferred tradeoff between coherence and attribution will vary depending on the application.
Space AmaSum
Clusters LLM R-2 $\uparrow$ R-L $\uparrow$ Prev. $\uparrow$ Gen. $\downarrow$ SAP $\uparrow$ R-2 $\uparrow$ R-L $\uparrow$ Prev. $\uparrow$ Gen. $\downarrow$ SAP $\uparrow$
Llama 2 7B Zero-shot Llama 2 7B 4.4 17.6 32.6 25.7 19.7 1.5 11.5 17.7 17.6 8.9
HIROsent Llama 2 7B 5.2 16.9 36.2 20.0 26.2 3.6 14.0 14.6 6.3 11.5
HIROdoc Llama 2 7B 6.9 19.9 48.5 29.9 33.5 4.0 15.2 16.0 12.1 10.0
Llama 2 13B Zero-shot Llama 2 13B 6.3 19.1 36.7 28.1 22.6 2.3 13.1 18.1 14.3 10.9
HIROsent Llama 2 13B 6.6 18.8 35.0 23.7 23.1 3.8 14.3 13.5 6.2 10.4
HIROdoc Llama 2 13B 8.5 21.7 46.2 27.0 32.7 4.3 15.7 19.0 9.9 14.1
Llama 2 70B Zero-shot Llama 2 70B 5.6 19.4 48.9 37.2 30.3 2.1 12.5 16.8 22.4 5.6
HIROsent Llama 2 70B 5.8 18.5 41.1 19.6 31.3 3.6 14.5 15.1 3.7 13.2
HIROdoc Llama 2 70B 8.3 21.3 48.5 30.0 33.5 4.4 16.0 17.8 9.9 12.9
Sent-wise $k$-means Mistral 7B 4.5 17.1 30.1 30.8 14.7 3.2 13.3 12.9 12.4 6.7
Hercules Mistral 7B 3.9 17.0 27.4 25.8 14.5 3.3 11.8 8.5 8.7 4.1
HIROsent Mistral 7B 4.5 18.2 36.4 20.2 26.3 3.5 14.1 14.6 6.9 11.2
Doc-wise $k$-means Mistral 7B 6.4 20.5 40.2 36.3 22.0 3.7 14.5 16.2 13.9 9.2
Hercules Mistral 7B 7.6 21.0 42.9 39.2 23.3 3.8 14.7 13.8 11.6 8.0
HIROdoc Mistral 7B 7.0 20.5 44.0 28.8 29.6 4.0 15.1 15.3 9.4 10.6
Table 4: Automatic evaluations comparing HIRO to zero-shot summarization, using a range of different LLMs, for the test splits. We show the mean scores based on 3 samples, and best scores within each comparison are bolded. In all cases, HIRO improves on the zero-shot approach. We also compare HIRO to other cluster selection methods, finding that HIRO leads to summaries with a better balance between prevalance and specificity.
@@
System Output
Reference The staff were very friendly, courteous and helpful. The hotel and room was clean. The rooms are very spacious with a nice oceanfront view and a gigantic balcony! Pool bar is great and food is good, too. The hotel has a perfect location. It was very nicely located with a 24 hours cafe downstairs, 2 pools , 2 bars and some beach stores to shop in.
Mistral 7B The Grand Hotel in Ocean City, Maryland, consistently receives positive reviews for its exceptional customer service, cleanliness, and convenient location on the boardwalk. Rooms offer great views, and amenities include a spa, restaurant, lounge, indoor pool, game room, and exercise room. Guests appreciate the family-friendly atmosphere and the ease of having everything they need within the hotel. Some reviewers praise Nina Railean for her helpfulness and accommodating nature. However, one review warns against staying during Memorial Day weekend due to loud parties.
HIROsent All rooms offer excellent ocean views at this hotel. The Grand Hotel & Spa offers direct ocean views and easy beach access. The Grand Hotel & Spa offers clean and spacious rooms. No problems with accommodations or access despite rain and floor issues. The Grand Hotel & Spa offers beautiful views and a comfortable stay with excellent service. The Grand Hotel & Spa exceeded expectations. Rooms need updating peeling wallpaper, large, dark interiors.
HIROdoc The Grand Hotel & Spa offers stunning ocean views from most rooms, with some having a direct view of the boardwalk and beach. The hotel’s design ensures that guests can enjoy the ocean despite room location. The beach is clean, and the location is convenient for easy access. Rooms are described as clean and large, but some note a need for improvement in terms of lighting and room condition. Overall, guests had positive experiences and appreciated the ocean access.
Table 5: Generated summaries for The Grand Hotel Maryland, from Space, comparing HIRO to zero-shot Mistral 7B. While all generated summaries are fluent and convincing, Mistral 7B makes reference to the opinion of a single user, and positive sentiment about a member of staff that is not supported by the full set of reviews. These extrapolations highlight the problem with models that can only accept a limited number of reviews as input.
HIRO provides more specific detail about user frustrations and the product itself. Mistral 7B is quite generic.
Mistral 7B mentions that desk staff were unfriendly, but this is not substantiated by the reviews, the majority of which are overwhelmingly positive. It’s also a contradiction since earlier it says the staff were friendly.
Only HIRO references the high failure rate.
HIRO seems more accurate in its appraisal of the staff; the reviews were mixed. However Mistral 7B is slightly more informative, particularly in relation to location and proximity to amenities. I feel like HIRO sits on the fence quite a lot, whereas Mistral 7B attempts to summarise better.
I think HIRO is the better written and reflects a broader view on the reviews but References isn’t bad it does cover PS4 and gaming which alot of reviews were using it for.
SemAE is incoherent. HIRO does an excellent job of summarising everything, balancing all perspectives.
Can tell HIRO is AI generated
HIRO is by far the better written and has a little more detail and reflects the reviews more than Herculesabs
Table 6: Comments from annotators for all pairs involving HIRO. Anonymised system labels have been replaced with the system names. The majority of comments are positive towards HIRO, although annotators comment that HIRO summaries may be less natural, or too conservative.
6.4 Analysis
Ablations
Table 4 compares HIRO to zero-shot prompting, for a range of different LLM sizes from the Llama 2 Chat family of LLMs (Touvron et al., 2023). The same prompt template and hyperparameters were used as for Mistral 7B. Increased parameter count does lead to improved SAP scores for the zero-shot approach on Space, but the largest 70B model scores worse on AmaSum than the 7B and 13B models. For all choices of LLM, HIRO leads to summaries with a better balance between prevalence and specificity than zero-shot prompting; HIRO is likely to lead to improvements for any choice of instruction-tuned LLM.
We also compare using HIRO to identify clusters against a $k$-means baseline, and Hercules (Hosking et al., 2023b). For both datasets, using clusters selected with HIRO leads to summaries that are much less generic, while remaining comparatively prevalent. This confirms the results in Section 5.2, indicating that HIRO selects more informative clusters of sentences than the comparison methods.
Qualitative Analysis
Table 5 shows output from Mistral 7B and HIRO as well as a reference summary, for The Grand Hotel Maryland from Space. While all generated summaries are fluent and convincing, Mistral 7B makes reference to the opinion of a single user, which should not appear in a summary. It also describes positive praise about a member of staff, but a manual analysis shows that only 3 out of 5 reviews mentioning that individual are positive; the true sentiment is much more mixed than the summary indicates. These extrapolations highlight the problem with models that can only accept a limited number of reviews as input.
The examples of selected subpaths, sentence clusters and corresponding HIRO outputs in LABEL:tab:evidence_example demonstrate the difficulty of evaluating attribution. The final cluster contains sentences that are all topically related, indicating that HIRO has learned a successful clustering. While the sentences and corresponding HIROsent output are all broadly negative, it is not straightforward to determine whether the sentence “Only one mirror in the room" counts as direct evidence towards the output “Rooms were small, loud, and in need of renovation […]". This partly explains the relatively low majority support scores in Table 3; some of the selected evidence may be consistent in topic and sentiment, but not directly entail the resulting output.
As well as collecting pairwise preferences in our human evaluation, we allowed annotators to leave qualitative comments. Table 6 shows all non-trivial comments from annotators for pairs including HIRO, with anonymised labels replaced by the true model names. The majority of comments are positive towards HIRO, highlighting improved levels of detail and a better balance of the input reviews. However, some annotators note that HIRO summaries may be less natural or too conservative.
@@ 7 Conclusion
We propose HIRO, a modular method for unsupervised opinion summarization that uses a hierarchical index over sentences to select clusters of prevalent opinions. Our approach leverages pretrained Large Language Models to generate coherent and fluent summaries that are attributable and accurately reflect the popularity of opinions in the input reviews. Extensive experiments show that, as well as generating higher quality summaries, HIRO learns a more semantically distributed representation than competitive baselines. While we limit our experiments to opinion summarization, we believe that HIRO could be usefully applied to a wide range of other retrieval-augmented generation tasks.
Acknowledgements
We thank the action editor and anonymous reviewers for their constructive feedback. This work was supported in part by the UKRI Centre for Doctoral Training in Natural Language Processing, funded by the UKRI (grant EP/S022481/1) and the University of Edinburgh. Lapata acknowledges the support of the UK Engineering and Physical Sciences Research Council (grant EP/W002876/1).
@@ Appendix B Hyperparameters
Optimizer Adam (Kingma and Ba, 2015)
@@ Appendix C LLM prompts
Prompt 1: Baseline LLMs (Mistral 7B and all Llama 2 models)